OpenStudy (anonymous):

Each of two identical springs will stretch 12 cm when a 250-g mass is hung on them. If the two springs are hung side by side and both support the 250-g mass, how much will each spring stretch?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

it sounds like each spring will support 1/2 the weight. Because they follow Hooke's Law, they will stretch 1/2 as much as when they have to support the whole weight.

OpenStudy (anonymous):

so the formula?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

The formula is written F= k x where F is the force , x is the distance stretched, and k is the "spring constant"

OpenStudy (anonymous):

F= 250*20

OpenStudy (anonymous):

@phi

OpenStudy (phi):

Force is mass times acceleration. in this case, it is the force due to gravity, and it is F= mass * g (g is 9.8 m/s^2) so you could say 250*g = k * 0.12 m and k= 250*g/0.12 next, if you use 2 springs, each will support one half of the force, so you have 125*g = k x using the k = 250*g/0.12 you can solve for x you should find x= 0.06 meters = 6 cm